Given a population consists of three numbers (2,5,9). Consider all possible samples of size 2 which can be drawn from the population. Find the standard deviation of the sampling distribution of the sample means. 


1
Expert's answer
2021-11-16T09:31:48-0500

possible samples mean

2,5 3.5

2,9 5.5

5,9 7


mean of the sample means:


"\mu=\frac{3.5+5.5+7}{3}=5.33"


standard deviation of the sample means:


"\sigma=\sqrt{\frac{\sum(\overline{x_i}-\mu)^2}{n}}=\sqrt{\frac{(3.5-5.33)^2+(5.5-5.33)^2+(7-5.33)^2}{3}}=1.43"